please help me solve this question:
Chloe is 2 years older than his husband damon. if the product of their ages is 30 more than 15 times the sum of their ages. how old is chloe
i know that this is simple but still i cant solve it

Answer by Maths68(1474)

Chole’s age = c
Husband’s age = h
Product of their ages = c*h
Sum of their ages = c+h
15 times the sum of their ages = 15 (c+h)
Chole two years older than her husband so if we add 2 years to the husband’s age both will be the same age
c=h+2 …….(1)
the product of their ages is thirty more than fifteen times the sum of their ages
c*h=30+15(c+h)……..(2)
Put the value of c from (1) to (2)
(h+2)*h=30+15[(h+2)+h]
h^2+2h=30+15h+30+15h
h^2+2h=60+30h
h^2+2h-60-30h=0
h^2-28h-60=0
h^2-30h+2h-60=0
h(h-30)-2(h-30)=0
(h+2)(h-30)
h=-2 h=30
h=-2 impossible
so h=30
Husband is 30 years old
And
Chole is 32